# DynamicVL: Benchmarking Multimodal Large Language Models for Dynamic City Understanding

### v1

@@ 4 DynamicVL Benchmark
Basic change analysis. The Sankey diagram in Fig. 6 quantifies land cover transition patterns between initial and final states. The visualization reveals a dominant trend of vegetation conversion to developed areas, reflecting rapid urbanization processes. In contrast, only a modest area of building, approximately $2.34km^{2}$ underwent demolition, primarily transitioning to vegetation and non-vegetation surfaces. These complex dynamics, involving multi-directional transitions among five distinct land cover types and requiring precise quantification across various spatial scales, present significant challenges for MLLMs in terms of both semantic understanding and numerical reasoning.
Change speed estimation. The temporal analysis in Fig. 6 tracks building expansion rates across successive periods, providing insights into U.S. urban development trajectories over the past two decades. Development velocity exhibits a distinct non-linear pattern: accelerating from 2010, reaching peak urbanization rates around 2017, and showing significant deceleration after 2018. This characteristic growth curve, with its pronounced acceleration and subsequent slowdown, represents a typical urban development cycle. Such complex temporal dynamics require MLLMs to maintain precise numerical sensitivity while modeling long-term spatiotemporal variations.
@@
Referring change detection. Analysis of change scales in Fig. 8 reveals distinct patterns across the five primary land-cover types. Changes in vegetation and non-vegetation areas show high variability, with a clear asymmetric distribution favoring smaller spatial extents. Buildings, water bodies, and recreational areas predominantly undergo small-scale changes, consistent with their limited spatial footprint in urban landscapes. These diverse change scales across land-cover types test MLLMs’ ability to detect and segment temporal changes at varying spatial scales.
@@
Figure 8: Word cloud in the dense temporal and regional-level change caption.
Dense temporal/Region-level change caption. The word cloud visualization in Fig. 8 illustrates the distribution of the 300 most frequent terms in our long-form captions. While existing geospatial vision-language datasets primarily focus on spatial descriptors (topright, left, etc.), DVL-Bench distinctively features rich temporal references (year, period, etc.) and transition vocabulary (replaced, added, completed, etc.), enabling comprehensive documentation of urban dynamics.
@@ 5 Benchmark Experiments
5.1 Implementation Details.
@@
Benchmark methods. We evaluate 17 widely-adopted MLLMs across different capabilities: (1) open-source MLLMs, including domain-specific models like TEOChat irvin2024teochat, state-of-the-art MLLMs with multi-image and video perception abilities (Video-LLaVA lin2024video, LLaVA-OneVision li2024llavaonevision, InternVL3 zhu2025internvl3, and Qwen2.5-VL bai2025qwen25vl), and referring segmentation models (LISA lai2024lisa, PSALM zhang2025psalm); (2) commercial MLLMs, including o4-mini openai2025o4mini, GPT4.1 openai2025gpt41, GPT4o openai2024gpt4o, and Gemini 2.5 Flash google2025gemini25flash. All experiments were conducted on 8 H100 GPUs. Unless otherwise specified, each model’s multi-image inference was employed for question-answering tasks, while LISA and PSALM were specifically evaluated on referring segmentation tasks. The detailed training and testing methods can be found in the Appendix §A.
@@
Evaluation metrics. Our evaluation framework employs multiple metrics to assess model performance across different tasks comprehensively. For the evaluations presented in Table 2, we measure both basic change analysis and change speed estimation using two approaches. First, we calculate accuracy percentages for single and multiple-choice questions. Second, for open-ended generation tasks, we evaluate Basic Change Reports using three metrics: Land Cover Type Identification (LCT), Time Period Accuracy (TPA), and Change Quantification Accuracy (CQA). Similarly, Change Speed Reports are assessed using Change Rate Precision (CRP), Time Period Accuracy (TPA), and Change Pattern Accuracy (CPA). For the long-form captioning tasks shown in Table 3, specifically regional-level change and dense temporal captioning, we evaluate Regional Change Captions using four metrics: Temporal Coverage (TC), Spatial Accuracy (SA), Process Fidelity (PF), and Region Containment (RC). All captioning metrics are scored on a 0-5 scale, with higher scores indicating better performance. These scores are determined by GPT-4.1-mini openai2025gpt41mini through comparison with reference captions. The detailed evaluation prompts can be found in the Appendix §B.
@@
Figure 9: Detailed illustration of DVLChat.We separate question-answering and referring change detection by implementing two distinct LoRA modules, enabling the model to possess independent VQA and segmentation capabilities while preventing interference between their respective data streams in the training.
@@
DVLChat design. As dynamic urban understanding necessitates both fine-grained pixel-level understanding and semantic comprehension, we followed the main architecture of LISA lai2024lisa to develop DVLChat. However, the original LISA model was unable to perform referring segmentation while maintaining high open-ended capabilities due to optimization conflicts. Furthermore, LISA, being designed for single-image analysis, lacked the ability to analyze changes for multiple images, where our DVL-Instruct data makes these capabilities possible. Leveraging our DVL-Instruct dataset, we develop DVLChat as a baseline model for multi-temporal urban understanding tasks. As shown in Figure 9, DVLChat employs a task-specific routing mechanism through dedicated prefix tokens from users. The system routes user queries to specialized modules based on their prefixes: inputs with [QA] activate the VQA LoRA module for generating textual responses, while those with [SE] engage the segmentation LoRA module. DVLChat addresses multi-temporal analysis by interleaving image features from multiple temporal images before decoding. For referring change detection tasks, the system processes this interleaved representation and decodes the <SEG> token embedding using SAM’s kirillov2023segment frozen vision backbone and unfrozen decoder to generate precise segmentation masks. DVLChat effectively isolates question-answering and segmentation functionalities, preventing task interference in the original LISA algorithm while maintaining model efficiency. While our implementation uses Qwen2.5-VL as the MLLM, the architecture is MLLM-agnostic and can accommodate other multimodal language models.
@@
LLaVA-OneVision 7B li2024llavaonevision 19.3 41.7 2.8 21.5 4.8 25.9 2.30 2.29 3.20 1.42 1.72 0.95 2.44 1.78
LLaVA-OneVision 72B li2024llavaonevision 25.0 59.9 6.5 25.9 6.2 26.5 3.01 2.70 4.52 1.83 2.05 0.93 3.39 1.83
@@
Qwen2.5-VL 3B bai2025qwen25vl 24.7 56.9 6.0 26.1 9.2 25.1 2.99 2.72 4.58 1.65 1.72 0.57 3.42 1.18
Qwen2.5-VL 7B bai2025qwen25vl 23.3 54.6 4.8 28.5 13.6 15.0 2.94 2.49 4.70 1.62 1.73 0.25 3.90 1.05
Qwen2.5-VL 32B bai2025qwen25vl 31.4 62.0 33.3 36.9 3.2 21.6 3.04 2.65 4.65 1.81 2.60 1.21 3.89 2.71
Qwen2.5-VL 72B bai2025qwen25vl 29.7 65.4 24.3 34.6 4.0 20.2 2.99 2.61 4.64 1.71 2.27 0.72 3.76 2.33
@@
Overall benchmark performance. As shown in Table 2, quantitative results reveal significant challenges in understanding urban dynamics through remote sensing imagery, with all current models demonstrating limited capabilities. The highest overall accuracy reaches merely 34.1% with o4-mini, while Qwen2.5-VL 32B and GPT4.1 achieve 31.4% and 32.5% respectively. Notably, TEOChat, despite being specifically designed for multi-temporal remote sensing vision-language tasks, achieves only 17.2% overall accuracy, struggling significantly with the benchmark’s larger and city-level understanding compared to its native $256\times 256$ input size. In contrast, our DVLChat, leveraging the newly proposed DVL-Instruct dataset, demonstrates competitive performance at 33.3% while maintaining high-resolution ($1024\times 1024$) processing capabilities.
@@
Task-specific challenges. Diverse tasks evaluate MLLMs’ performances from different aspects. While [BCA] with single-choice questions shows promising results, where Qwen2.5-VL 72B achieves 65.4% accuracy, performance degrades substantially in multi-choice settings, where even the leading model GPT4.1 only reaches 39.7%. The challenges become more pronounced in detailed analytical tasks. [BCA] report metrics expose fundamental limitations in both land cover type identification and change quantification, with LCT scores maxing at 2.85 and CQA not exceeding 1.93 across existing models. Notably, by leveraging DVL-Instruct’s comprehensive training data, DVLChat achieves a significant breakthrough in LCT with a score of 3.41, enhancing the recognition of changed land-cover types. [CSE] results reveal a critical limitation of current MLLMs in pixel-level change perception, with multi-choice accuracy peaking at merely 13.6% and CRP consistently below 1.21, indicating models’ inability to capture and quantify fine-grained temporal variations. [EA] results are similarly concerning, with accuracies ranging from 14.6% to 26.8%, where many models perform close to or even below random chance (20% for 5-options).
@@
Captioning capabilities. Table 3 reveals a substantial gap between commercial and open-source models in regional change captioning tasks. Commercial models demonstrate superior performance with o4-mini achieving an average score of 4.58, while the best open-source model, InternVL3 14B, reaches only 3.96. Our DVLChat, through the incorporation of DVL-Instruct, demonstrates strong performance with an average score of 3.98, approaching commercial model performance. The disparity becomes even more pronounced in dense temporal captioning, where commercial models maintain strong performance with o4-mini reaching an average score of 4.14, while open-source alternatives struggle considerably with scores below 3.40. Notably, TEOChat achieves only 1.45, revealing severe limitations in handling complex temporal dynamics beyond bi-temporal comparisons.
@@
Table 3: Performance comparison of different models on regional change caption and dense temporal caption tasks, evaluated using Temporal Coverage (TC), Spatial Accuracy (SA), Process Fidelity (PF), and Region Containment (RC) metrics on a 0-5 scale.
Method Regional Change Caption Dense Temporal Caption
@@
LLaVA-OneVision 7B li2024llavaonevision 3.07 3.12 2.37 2.17 4.63 2.17 2.36 2.09 2.08
LLaVA-OneVision 72B li2024llavaonevision 3.60 3.91 2.77 2.80 4.93 2.87 3.51 2.56 2.54
@@
Qwen2.5-VL 3B bai2025qwen25vl 2.76 2.77 1.82 1.52 4.92 2.38 2.38 2.66 2.11
Qwen2.5-VL 7B bai2025qwen25vl 3.21 3.30 2.42 2.20 4.92 2.85 3.47 2.57 2.51
Qwen2.5-VL 32B bai2025qwen25vl 3.90 4.28 3.18 3.23 4.92 2.91 3.39 2.77 2.57
Qwen2.5-VL 72B bai2025qwen25vl 3.89 4.24 3.24 3.17 4.90 3.28 3.94 2.95 2.95
@@
Scaling parameters. Analysis of model size scaling reveals inconsistent improvements within different model families, which is distinguished from most generic computer vision tasks yue2024mmmupro; hu2025videommmu. Particularly in basic change analysis and change speed estimation tasks, the Qwen2.5-VL series shows notable improvements as model size increases to 32B, reaching 31.4% overall accuracy, but performance declines to 29.7% with the 72B counterpart. Similarly, while LLaVA-OneVision improves from 19.3% to 25.0% when scaling from 7B to 72B, InternVL3 peaks at 27.2% with its 14B variant before slightly declining to 27.1% with the 78B model. These non-monotonic scaling patterns in analytical tasks contrast sharply with the consistent improvements observed in captioning tasks, where larger models consistently achieve better performance in both regional and dense temporal captioning. This divergence in scaling behavior suggests that while model size benefits language generation and temporal narrative abilities, merely increasing parameters is insufficient for enhancing precise change detection and quantification capabilities. Instead, domain-specific data incorporating strategies may be crucial for advancing model capabilities in these analytical aspects of urban dynamics understanding.
Referring change detection analysis. We fine-tuned the state-of-the-art specialist change detection model (ChangeMamba chen2024changemamba) and the MLLM-based referring segmentation models (LISA lai2024lisa and PSALM zhang2025psalm) for comparison. As shown in Fig. 10, the specialist ChangeMamba achieves high performance by training on fixed labels. Regarding MLLMs, LISA demonstrates superior F1 relative to PSALM, but achieves lower IoU. Because foreground $\ll$ background pixels in change detection scenarios, even if most pixels are classified correctly (improving F1), the predicted area may have a low overlap with the ground-truth area (low IoU). Equipped with a stronger backbone MLLM, i.e., Qwen2.5-VL, our DVLChat achieves optimal performance, further narrowing the gap compared to the specialist model and providing a reliable baseline for future research.
@@
Limitations. Several key directions remain for future exploration. First, DVL-Suite includes near-infrared band information from NAIP imagery, but current MLLMs’ limited capabilities in processing these spectral bands prevent their potential utilization, particularly for economic assessment tasks. Second, while DVLChat provides a unified baseline, it does not yet leverage pixel-level segmentation data to enhance numerical quantification across tasks. Finally, while DVLChat outperforms existing open-source models on most metrics, it still falls behind commercial models. Future work will focus on developing specialized algorithms and scaling up parameter size to bridge this performance gap.

### v2

@@ 4 DVL-Bench Designs
Basic change analysis (BCA). The Sankey diagram in Figure 6 quantifies land cover transition patterns between initial and final states. The visualization reveals a dominant trend of vegetation conversion to developed areas, reflecting rapid urbanization processes. In contrast, only a modest area of building, approximately $2.34km^{2}$ underwent demolition, primarily transitioning to vegetation and non-vegetation surfaces. These complex dynamics, involving multi-directional transitions among five distinct land cover types and requiring precise quantification across various spatial scales, present significant challenges for MLLMs in terms of both semantic understanding and numerical reasoning.
Change speed estimation (CSE). The temporal analysis in Figure 6 tracks building expansion rates across successive periods, providing insights into U.S. urban development trajectories over the past two decades. Development velocity exhibits a distinct non-linear pattern: accelerating from 2010, reaching peak urbanization rates around 2017, and showing significant deceleration after 2018. This characteristic growth curve, with its pronounced acceleration and subsequent slowdown, represents a typical urban development cycle. Such complex temporal dynamics require MLLMs to maintain precise numerical sensitivity while modeling long-term spatiotemporal variations.
@@
Referring change detection (RCD). Analysis of change scales in Figure 8 reveals distinct patterns across the five primary land-cover types. Changes in vegetation and non-vegetation areas show high variability, with a clear asymmetric distribution favoring smaller spatial extents. Buildings, water bodies, and recreational areas predominantly undergo small-scale changes, consistent with their limited spatial footprint in urban landscapes. These diverse change scales across land-cover types test MLLMs’ ability to detect and segment temporal changes at varying spatial scales.
@@
Figure 8: Word cloud in the dense temporal and regional change captioning tasks.
Dense temporal and region-level change caption (DTC and RCC). The word cloud visualization in Figure 8 illustrates the distribution of the 300 most frequent terms in designed long-form captions. While existing geospatial vision-language datasets primarily focus on spatial descriptors (topright, left, etc.), DVL-Bench distinctively features rich temporal references (year, period, etc.) and transition vocabulary (replaced, added, completed, etc.), enabling comprehensive documentation of urban dynamics.
@@ 5 Benchmark Experiments
In this section, we present the experimental setup, introduce DVLChat as our baseline model for dynamic urban understanding, and comprehensively evaluate 18 state-of-the-art MLLMs on DVL-Bench.
5.1 Implementation Details
@@
Benchmark methods. We evaluate 18 widely-adopted MLLMs across different capabilities: (1) open-source MLLMs, including domain-specific models like TEOChat irvin2024teochat and EarthDial soni2024earthdial, state-of-the-art MLLMs with multi-image and video perception abilities (Video-LLaVA lin2024video, LLaVA-OneVision li2024llava-ov, InternVL3 zhu2025internvl3, and Qwen2.5-VL bai2025qwen25), and referring segmentation models (LISA lai2024lisa, PSALM zhang2025psalm); (2) commercial MLLMs, including o4-mini openai2025o4mini, GPT4.1 openai2025gpt41, GPT4o openai2024gpt4o, and Gemini 2.5 Flash google2025gemini25flash. Unless otherwise specified, all experiments using open-source models were conducted on 8 H100 GPUs. For question-answering tasks, we utilized each model’s native multi-image inference capability. For referring change detection tasks, LISA and PSALM were evaluated using a different approach, concatenating two temporal images into a single composite image as input. The detailed training and testing methods can be found in the Appendix § A.
@@
Evaluation metrics. Our evaluation framework employs multiple metrics to assess model performance across different tasks comprehensively. For the evaluations presented in Table 2, we measure both basic change analysis and change speed estimation using two approaches. First, we calculate accuracy percentages for single and multiple-choice questions. Second, for open-ended generation tasks, we evaluate Basic Change Reports using three metrics: Land Cover Type Identification (LCT), Time Period Accuracy (TPA), and Change Quantification Accuracy (CQA). Similarly, Change Speed Reports are assessed using Change Rate Precision (CRP), Time Period Accuracy (TPA), and Change Pattern Accuracy (CPA). For the long-form captioning tasks shown in Table 3, Regional Change Captioning is evaluated using Temporal Coverage (TC), Spatial Accuracy (SA), Process Fidelity (PF), and Region Containment (RC), while Dense Temporal Captioning uses TC, SA, and PF. All captioning metrics are scored on a 0-5 scale, with higher scores indicating better performance. These scores are determined by GPT4.1-mini openai2025gpt41mini through comparison with reference captions. The detailed evaluation prompts can be found in the Appendix § B.2.
@@
Figure 9: Detailed illustration of DVLChat.We separate question-answering and referring change detection by implementing two distinct LoRA hu2022lora modules, enabling the model to possess independent VQA and segmentation capabilities while preventing interference between their respective data streams in the training.
@@
DVLChat design. As dynamic urban understanding necessitates both semantic comprehension and fine-grained pixel-level understanding, we followed the main architecture of LISA lai2024lisa to develop DVLChat. However, the original LISA model was unable to perform pixel-level segmentation while maintaining high open-ended capabilities due to optimization conflicts. Furthermore, LISA, designed for single-image analysis, lacked the ability to analyze changes across multiple images, whereas the DVL-Instruct data enables these capabilities. Therefore, leveraging DVL-Instruct, we develop DVLChat as a baseline model for multi-temporal urban understanding tasks. As shown in Figure 9, DVLChat employs a task-specific routing mechanism through dedicated prefix tokens from users. The system routes user queries to specialized modules based on their prefixes: inputs with [QA] activate the VQA LoRA module for generating textual responses, while those with [SE] engage the change detection LoRA module. DVLChat addresses multi-temporal analysis by interleaving image features from multiple temporal images before decoding. For referring change detection tasks, the system processes this interleaved representation and decodes the <SEG> token embedding using SAM’s kirillov2023segment frozen vision backbone and unfrozen decoder to generate precise segmentation masks. DVLChat effectively isolates question-answering and change detection functionalities, preventing task interference in the original LISA algorithm while maintaining model efficiency. While our implementation uses Qwen2.5-VL as the MLLM, the architecture is MLLM-agnostic and can accommodate other multimodal language models.
@@
LLaVA-OneVision 7B li2024llava-ov 19.3 41.7 2.8 21.5 4.8 25.9 2.30 2.29 3.20 1.42 1.72 0.95 2.44 1.78
LLaVA-OneVision 72B li2024llava-ov 25.0 59.9 6.5 25.9 6.2 26.5 3.01 2.70 4.52 1.83 2.05 0.93 3.39 1.83
@@
Qwen2.5-VL 3B bai2025qwen25 24.7 56.9 6.0 26.1 9.2 25.1 2.99 2.72 4.58 1.65 1.72 0.57 3.42 1.18
Qwen2.5-VL 7B bai2025qwen25 23.3 54.6 4.8 28.5 13.6 15.0 2.94 2.49 4.70 1.62 1.73 0.25 3.90 1.05
Qwen2.5-VL 32B bai2025qwen25 31.4 62.0 33.3 36.9 3.2 21.6 3.04 2.65 4.65 1.81 2.60 1.21 3.89 2.71
Qwen2.5-VL 72B bai2025qwen25 29.7 65.4 24.3 34.6 4.0 20.2 2.99 2.61 4.64 1.71 2.27 0.72 3.76 2.33
@@ 5.2 Benchmark Results
Overall benchmark performance. As shown in Table 2, quantitative results reveal significant challenges in understanding urban dynamics through remote sensing imagery, with all current models demonstrating limited capabilities. The highest averaged accuracy of multiple-choice questions reaches merely 34.1% with o4-mini, while Qwen2.5-VL 32B and GPT4.1 achieve 31.4% and 32.5% respectively. Notably, TEOChat, despite being specifically designed for multi-temporal remote sensing vision-language tasks, achieves only 17.2% overall accuracy, struggling significantly with the benchmark’s larger and city-level understanding compared to its native $256\times 256$ input size. In contrast, our DVLChat 7B, leveraging the proposed DVL-Instruct dataset, demonstrates competitive performance at 33.3% while maintaining referring change detection capabilities.
@@
Task-specific challenges. Diverse tasks evaluate MLLMs’ performances from different aspects. While [BCA] with single-choice questions shows promising results, where Qwen2.5-VL 72B achieves 65.4% accuracy, performance degrades substantially in multi-choice settings, where even the leading model GPT4.1 only reaches 39.7%. The challenges become more pronounced in detailed analytical tasks. [BCA] report metrics expose fundamental limitations in both land cover type identification (LCT) and change quantification (CQA), with LCT scores maxing at 2.85 and CQA not exceeding 1.93 across existing models. Notably, by leveraging DVL-Instruct’s comprehensive training data, DVLChat achieves a significant breakthrough in LCT with a score of 3.41, enhancing the recognition of changed land-cover types. [CSE] results reveal a critical limitation of current MLLMs in pixel-level change perception, with multi-choice accuracy peaking at merely 13.6% and Change Rate Precision (CRP) consistently below 1.21, indicating models’ inability to capture and quantify fine-grained temporal variations. [EA] results are similarly concerning: except for our DVLChat 7B, other models achieve accuracies ranging from 14.6% to 26.8%, with many performing at or even below random chance (20% for 5-option questions).
@@
Captioning capabilities. Table 3 reveals a substantial gap between commercial and open-source models in detailed captioning tasks. For the regional change captioning task, commercial models demonstrate superior performance with o4-mini achieving an average score of 4.58, while the best open-source model, InternVL3 14B, reaches only 3.96. Our DVLChat, incorporating DVL-Instruct, demonstrates strong performance with an average score of 3.98, approaching the performance of commercial models with 7B parameters. The disparity becomes even more pronounced in dense temporal captioning, where commercial models maintain strong performance with o4-mini reaching an average score of 4.14, while open-source alternatives struggle considerably with scores below 3.40. Notably, TEOChat achieves only 1.45, revealing severe limitations in handling complex temporal dynamics beyond bi-temporal comparisons.
@@
Table 3: Performance comparison of different models on regional change captioning (RCC) and dense temporal captioning (DTC) tasks, evaluated using Temporal Coverage (TC), Spatial Accuracy (SA), Process Fidelity (PF), and Region Containment (RC) metrics on a 0-5 scale.
Method RCC DTC
@@
LLaVA-OneVision 7B li2024llava-ov 3.07 3.12 2.37 2.17 4.63 2.17 2.36 2.09 2.08
LLaVA-OneVision 72B li2024llava-ov 3.60 3.91 2.77 2.80 4.93 2.87 3.51 2.56 2.54
@@
Qwen2.5-VL 3B bai2025qwen25 2.76 2.77 1.82 1.52 4.92 2.38 2.38 2.66 2.11
Qwen2.5-VL 7B bai2025qwen25 3.21 3.30 2.42 2.20 4.92 2.85 3.47 2.57 2.51
Qwen2.5-VL 32B bai2025qwen25 3.90 4.28 3.18 3.23 4.92 2.91 3.39 2.77 2.57
Qwen2.5-VL 72B bai2025qwen25 3.89 4.24 3.24 3.17 4.90 3.28 3.94 2.95 2.95
@@
Scaling parameters. Analysis of model size scaling reveals inconsistent improvements within different model families, which is distinguished from most generic computer vision tasks yue2024mmmupro; hu2025videommmu. Particularly in basic change analysis and change speed estimation tasks, the Qwen2.5-VL series shows notable improvements as model size increases to 32B, reaching 31.4% average accuracy, but performance declines to 29.7% with the 72B counterpart. Similarly, while LLaVA-OneVision improves from 19.3% to 25.0% when scaling from 7B to 72B, InternVL3 peaks at 27.2% with its 14B variant before slightly declining to 27.1% with the 78B model. These non-monotonic scaling patterns in analytical tasks contrast sharply with the consistent improvements observed in captioning tasks, where larger models consistently achieve better performance in both regional and dense temporal captioning. This divergence in scaling behavior suggests that while model size benefits language generation and temporal narrative abilities, merely increasing parameters is insufficient for enhancing precise change detection and quantification capabilities. This is further evidenced by our DVLChat 7B outperforming larger models (up to 78B parameters) across multiple tasks when trained with domain-specific data. This highlights that incorporating strategies into domain-specific data is crucial for advancing model capabilities in understanding the analytical aspects of urban dynamics. We provide more analysis on scaling patterns by incorporating domain-specific data in Appendix § E.
Referring change detection analysis. We compare DVLChat with the specialist change-detection model ChangeMamba chen2024changemamba and MLLM-based referring-segmentation models LISA lai2024lisa and PSALM zhang2025psalm, all fine-tuned on our dataset. As shown in Figure 10, ChangeMamba attains the highest IoU (32.41%) as a task-specific model trained on a fixed target ("new buildings"). Among MLLM-based methods, PSALM (26.93%) outperforms LISA (13.85%). DVLChat reaches 29.06% IoU, within 3.35% of the specialist. The qualitative results show DVLChat’s tighter alignment with the ground truth than LISA/PSALM, especially around building boundaries and the spatial extent of new constructions.
@@
6 Limitations and Future Directions
Several key directions remain for future exploration. First, DVL-Suite includes near-infrared band information from NAIP imagery, but the limited capabilities of current MLLMs in processing these spectral bands prevent their potential utilization, particularly for economic assessment tasks. Second, while DVLChat provides a unified baseline, it does not yet leverage pixel-level segmentation data to enhance numerical quantification across tasks. Finally, while DVLChat outperforms existing open-source models on most metrics, it still falls behind commercial models. Future work will focus on developing specialized algorithms and scaling up parameter size to bridge this performance gap.